### v1

# Contextuality from Single-State Representations: An Information-Theoretic Principle for Adaptive Intelligence

Abstract
Adaptive systems often operate across multiple contexts while reusing a fixed internal state space due to constraints on memory, representation, or physical resources. Such single-state reuse is ubiquitous in natural and artificial intelligence, yet its fundamental representational consequences remain poorly understood. We show that contextuality is not a peculiarity of quantum mechanics, but an inevitable consequence of single-state reuse in classical probabilistic representations. Modeling contexts as interventions acting on a shared internal state, we prove that any classical model reproducing contextual outcome statistics must incur an irreducible information-theoretic cost: dependence on context cannot be mediated solely through the internal state. We provide a minimal constructive example that explicitly realizes this cost and clarifies its operational meaning. We further explain how nonclassical probabilistic frameworks avoid this obstruction by relaxing the assumption of a single global joint probability space, without invoking quantum dynamics or Hilbert space structure. Our results identify contextuality as a general representational constraint on adaptive intelligence, independent of physical implementation.
I Introduction
Adaptive systems—biological or artificial—must routinely operate across multiple contexts while relying on limited internal resources. In many realistic settings, such systems are constrained to maintain a single internal state space that is reused across different interactions, measurements, or decision scenarios. Here, “single-state” does not refer to a scalar or binary variable, but to a fixed internal state space—potentially high-dimensional and carrying many bits of information—that cannot be duplicated, indexed, or branched according to context. Such constraints arise naturally in perceptual systems with limited working memory, decision-making agents with bounded internal representations, and physical devices whose internal degrees of freedom must be reused across contexts. Related constraints have long been discussed across cognitive science and artificial intelligence under notions such as bounded rationality, resource-limited agents, and fixed internal representational architectures (e.g., Herbert A. Simon; Models of Bounded Rationality [4]; John R. Anderson [1]). Despite their ubiquity, the representational consequences of maintaining a fixed internal state space across contexts remain poorly understood.
A central difficulty lies in how contextual dependence is represented under such constraints. Standard classical approaches typically address contextuality by introducing external labels, hidden variables, or context-dependent internal states, effectively enlarging the representational space. While successful in many applications, these strategies implicitly relax the single-state constraint by allowing contextual information to be stored or accessed explicitly. In contrast, many adaptive systems—such as perceptual agents responding to sequential stimuli under limited working memory or decision systems operating under strict memory reuse—must accommodate contextual changes through interventions that modify information processing, rather than through explicit encoding of which context is active (e.g., Working Memory [2]). This distinction motivates a fundamental question: what representational and informational costs are unavoidable when a system is required to maintain a fixed internal state space across multiple interventions?
In this work, we address this question at the level of representation rather than dynamics or physical implementation. We consider systems whose internal state space may be high-dimensional, but must remain fixed and internally consistent across contexts. Within this minimal framework, we show that maintaining such representations necessarily entails an irreducible information-theoretic cost when contextual dependence is present. Crucially, this cost does not arise from insufficient state capacity, but from the requirement that a single internal state space be reused coherently across multiple interventions. This perspective reframes contextuality not as a peculiarity of particular physical theories, but as a general and unavoidable constraint on representation.
Our approach makes three complementary contributions. First, we introduce a minimal operational formulation of fixed-state representations that cleanly separates contextual effects from assumptions about dynamics or physical realization. Second, we demonstrate that any classical probabilistic representation satisfying this formulation necessarily incurs an irreducible information-theoretic cost when contextual dependence is present, reflecting a fundamental obstruction to representing multiple interventions within a single internally consistent state space. Third, we present a simple constructive mechanism showing how nonclassical statistical structure can emerge directly from representational constraints, without invoking quantum postulates or Hilbert space assumptions. Together, these results establish contextuality as a representational principle, rather than a feature tied to specific physical theories.
The remainder of this paper is organized as follows. In Sec. II, we introduce a representational framework based on single-state reuse and contextual interventions, and define the class of classical probabilistic representations considered throughout. In Sec. III, we establish the central information-theoretic obstruction, proving that classical single-state representations necessarily incur an irreducible contextual information cost under contextual dependence. Sec. IV provides a minimal constructive illustration showing that this obstruction corresponds to realizable representational behavior rather than a purely formal limitation. In Sec. V, we explain why nonclassical probabilistic frameworks naturally accommodate single-state reuse without additional information cost. Sec. VI discusses the implications of these results for adaptive and intelligent systems operating under resource constraints, and Sec. VII concludes with a discussion of limitations and future directions.
II Representational Framework
II.1 Single-State Representations and Interventions
We begin by formalizing the representational framework considered throughout this work. Our goal is to isolate constraints on representation itself, independently of dynamics, learning rules, or physical implementation.
Definition 1 (Single-State Representation).
A single-state representation consists of a fixed internal state space $\mathcal{S}$ together with a family of response functions that map elements of $\mathcal{S}$ to observable outcomes. The state space $\mathcal{S}$ may be high-dimensional and encode many bits of information, but it is not indexed, duplicated, or partitioned according to context. All contextual variation must therefore be mediated through transformations acting on $\mathcal{S}$, rather than through context-dependent enlargement or branching of the state space.
Importantly, the single-state condition does not impose a bound on state capacity. Instead, it constrains how internal information may be reused across contexts: the same underlying state space must support all contextual interactions.
Definition 2 (Interventions).
A context is modeled as an intervention, defined as a map that acts on the internal state space $\mathcal{S}$ or on the response functions associated with $\mathcal{S}$. Interventions modify how the internal state is processed or probed, but do not introduce additional state variables or external context labels accessible to the representation. In particular, the identity of the intervention is not assumed to be explicitly encoded within $\mathcal{S}$.
This notion distinguishes interventions from classical context labels: while labels index distinct representational resources, interventions act on a common resource that must be reused across contexts.
Definition 3 (Representation Consistency).
A single-state representation is said to be consistent if all interventions act on the same internal state space $\mathcal{S}$ and yield outcome statistics that are jointly representable without contradiction. Consistency requires that the internal representation does not depend on which interventions are applied before or after, except through their defined action on $\mathcal{S}$. In particular, no intervention-specific hidden states or context-dependent refinements of $\mathcal{S}$ are permitted.
This condition enforces that the internal state retains a unified semantic interpretation across all interventions, rather than serving as a collection of unrelated context-specific representations.
Remarks.
Together, these definitions capture a minimal but stringent notion of representation reuse. They allow for rich internal structure while prohibiting the use of contextual indexing as a representational shortcut. The remainder of this work explores the consequences of these constraints, showing that classical probabilistic models cannot, in general, satisfy them without incurring an irreducible information-theoretic cost.
II.2 Classical Probabilistic Representations
We now specify what we mean by a classical probabilistic representation within the single-state framework introduced above. Throughout, we adopt a representation-level notion of classicality, independent of any specific physical implementation.
Definition 4 (Classical Probabilistic Representation).
A classical probabilistic representation is a single-state representation in which the internal state $s\in\mathcal{S}$ is associated with a classical random variable, and outcome statistics are generated by conditional probability distributions over observable outcomes given $s$. All randomness is assumed to be classical, and joint statistics across interventions are assumed to admit a single underlying probability space.
This definition captures the standard notion of classical simulation: observable statistics arise from probabilistic mixtures over underlying internal states, with no additional contextual degrees of freedom beyond those encoded in $\mathcal{S}$.
Crucially, under the single-state constraint, the same internal state space $\mathcal{S}$ must support all interventions without contextual indexing. Any dependence on context must therefore be mediated through the action of interventions on $\mathcal{S}$, rather than through context-dependent refinement of the underlying probability space.
Definition 5 (Context Information).
We define the context information as the amount of information required to account for contextual dependence in observable statistics beyond what can be encoded in the internal state $s\in\mathcal{S}$. This quantity measures the additional representational resources needed to reconcile multiple interventions within a single classical probability space.
If a classical probabilistic representation could faithfully encode all contextual dependence within the internal state alone, the required context information would vanish. However, as we show below, this is generically impossible under the single-state constraint.
The central question addressed in the remainder of this work can now be stated precisely: Can a classical probabilistic representation, constrained to a fixed internal state space reused across interventions, account for contextual dependence without incurring additional information-theoretic cost? In the next section, we show that the answer is negative. Specifically, we prove that any such representation necessarily requires an irreducible excess of contextual information, establishing a fundamental obstruction to classical single-state representations.
Operational random variables. To make information-theoretic statements explicit, we introduce the following random variables. Let $C$ denote the intervention (context) selected from a set $\mathcal{C}$, $S\in\mathcal{S}$ the internal state, and $O$ the observable outcome. A classical single-state model specifies conditional distributions $p(o\mid s,c)$ together with a prior over $S$. To represent contextual dependence without indexing $\mathcal{S}$ by $c$, the model may require additional contextual information $M$ (an external label or memory register) that is not contained in $S$. All entropies and mutual informations are Shannon quantities [3].
III Information-Theoretic Obstruction
III.1 Main Theorem
We now state the central result of this work. It establishes an unavoidable information-theoretic obstruction faced by classical probabilistic representations that maintain a fixed internal state space across multiple interventions.
Theorem 1 (Information-Theoretic Obstruction to Single-State Classical Representations).
Let $C$ denote the intervention, $S$ the internal state, and $O$ the outcome. Consider any classical single-state representation that reproduces the observed family of conditionals $\{p(o\mid c)\}_{c\in\mathcal{C}}$ via a single underlying probability space. If the observed statistics exhibit contextual dependence across interventions, then any such representation requires an additional contextual variable $M$ (not contained in $S$) such that
$H(M)\;\geq\;I(C;O\mid S)\;>\;0.$ (1)
Equivalently, the required contextual information strictly exceeds what can be attributed to the internal state alone. We write this irreducible excess in the shorthand form $H>I$ in the main text.
Operationally, $H(M)$ quantifies the additional contextual information (in Shannon bits, on average) required to reconcile contextual outcome statistics while preserving a single internal state.
III.2 Interpretation and Scope
The inequality in Eq. (1) is not remarkable by its mathematical form. Its significance lies in what it quantifies: an unavoidable representational cost arising from the requirement that a single internal state space be reused coherently across multiple interventions. Importantly, this obstruction does not reflect insufficient state capacity. Even arbitrarily large state spaces cannot eliminate the excess as long as contextual dependence must be supported within a single classical probability space.
The theorem is representation-theoretic and device-independent. It does not assume any particular dynamics, learning procedure, or physical substrate. Rather, it isolates a structural incompatibility between classical probabilistic representations and the requirement of single-state reuse in the presence of contextual dependence.
The proof proceeds by showing that, under the single-state constraint, contextual dependence forces incompatible conditional structures to be represented within a single underlying probability space. Resolving these incompatibilities requires additional information that cannot be encoded in the internal state without introducing context-dependent refinement. A full proof is provided in Appendix A.
IV A Constructive Illustration
The information-theoretic obstruction established in Sec. III is general and representation-theoretic. To show that this obstruction is not merely formal, we briefly introduce a minimal constructive illustration in which contextual dependence arises under the single-state constraint. The purpose of this illustration is purely explanatory: it serves as an existence witness demonstrating that the obstruction corresponds to realizable representational behavior, rather than to an abstract impossibility detached from concrete models.
$c_{1}$$c_{2}$$c_{3}$$p(o_{1},o_{2})$$p(o_{2},o_{3})$$p(o_{3},o_{1})$$\nexists\;p(o_{1},o_{2},o_{3})$
Figure 1: Schematic illustration of contextuality under the single-state constraint. Each pair of interventions admits a consistent marginal distribution, yet no global joint distribution exists that reproduces all pairwise marginals simultaneously. This incompatibility forces any classical single-state representation to introduce additional contextual information.
We consider a system with a fixed internal state space $\mathcal{S}$ that is reused across multiple interventions. Each intervention acts on the same internal state space and produces observable outcomes through intervention-dependent response functions, in accordance with the single-state and consistency conditions introduced in Sec. II. Crucially, no intervention identity is explicitly encoded within $\mathcal{S}$, and no context-dependent refinement or branching of the state space is permitted. Contextual variation is implemented solely through how interventions probe or transform the shared internal state.
As a minimal illustration, consider three interventions $c\in\{1,2,3\}$ with binary outcomes $o\in\{0,1\}$ such that each pair of interventions admits a consistent marginal distribution, while no single joint distribution over all three outcomes exists. Under the single-state constraint, reproducing these statistics requires additional contextual information specifying which incompatible marginal structure is realized, thereby instantiating the information-theoretic obstruction in a concrete setting.
For instance, let $p(o_{i},o_{j}\mid c_{i},c_{j})$ denote consistent pairwise marginals for each $(i,j)$, while no joint distribution $p(o_{1},o_{2},o_{3})$ exists that reproduces all marginals simultaneously. Any classical realization preserving a single internal state must therefore encode which marginal structure is active, requiring additional contextual information. Such marginal compatibility without global consistency is a standard signature of contextuality.
To make this explicit, consider a deterministic internal state $S\in\{0,1\}$ with $P(S=0)=P(S=1)=1/2$. Let the outcome be given by
$O=S\oplus f(C),$
where $f(C)\in\{0,1\}$ is an intervention-dependent bit. Here we assume that $S$ and $C$ are statistically independent, so that any dependence of $O$ on $C$ is not mediated through the internal state. Then $I(S;O)=0$, while
$I(C;O\mid S)=H(f(C))>0,$
since knowledge of $C$ is required to predict $O$ even when $S$ is known. This explicitly demonstrates that contextual dependence enforces $I(C;O\mid S)>0$, rendering the bound in Theorem 1 nontrivial.
Within this setting, one finds that although each intervention admits a well-defined probabilistic description in isolation, the resulting statistics across interventions cannot, in general, be embedded into a single classical probabilistic representation that preserves single-state reuse. The incompatibility does not arise from insufficient state capacity or from specific dynamical assumptions, but from the requirement that a single internal representation support multiple, mutually incompatible statistical dependencies. As a result, the observable behavior exhibits nonclassical statistical structure directly traceable to representational constraints.
This constructive behavior aligns with the general obstruction proven in Sec. III. Any attempt to classically simulate such behavior while maintaining a fixed internal state space necessarily incurs additional contextual information, consistent with the inequality $H>I$. Importantly, the illustration introduces no assumptions beyond those already formalized in the representational framework: it neither presupposes quantum dynamics nor invokes Hilbert space structure. Its sole role is to clarify how contextual dependence can arise operationally under single-state constraints and to provide intuition for the general information-theoretic result.
V Why Nonclassical Probability Appears
Nonclassical probabilistic frameworks circumvent this limitation by relaxing the requirement of a single global probability space. Instead, they allow probabilistic assignments that remain consistent within each intervention while preserving coherence at the level of the internal state. This structural flexibility enables contextual dependence to be represented without introducing external labels or hidden contextual memory, thereby avoiding the information-theoretic excess quantified by the inequality $H>I$.
From a representational perspective, the obstruction arises from the requirement that all observable statistics be embedded into a single global probability space. Nonclassical probabilistic frameworks relax precisely this requirement: probabilities are assigned consistently within each intervention, while global joint representability is replaced by a weaker consistency condition tied to the shared state description. This structural relaxation is sufficient to avoid encoding intervention identity as an explicit variable, without committing to any particular physical interpretation.
In a quantum representation, the state $\rho$ is single and fixed, but outcome probabilities are given by $p(o\mid c)=\mathrm{Tr}(\rho E_{o}^{(c)})$. Crucially, no joint probability distribution over outcomes for different interventions is assumed to exist. As a result, the classical requirement that all conditionals $p(o\mid c)$ arise from a single underlying random variable $S$ is absent. The quantity $I(C;O\mid S)$ is therefore not operationally meaningful in the same way, and no auxiliary contextual variable $M$ is required. From this perspective, quantum probability avoids the obstruction not by encoding context more efficiently, but by relaxing the assumption of global joint representability that underlies classical single-state models.
This viewpoint explains why nonclassical probability appears repeatedly in systems exhibiting contextual dependence, bounded memory, and order effects. Rather than signaling a departure from classical realism or physical intuition, it reflects a principled adjustment of representational structure necessitated by fixed internal state space across interventions.
VI Implications for Adaptive Intelligence
The results of this work have direct implications for adaptive and intelligent systems operating under internal resource constraints. Many such systems must interact with their environment across multiple contexts while reusing a fixed internal state space, due to limitations on memory, representational capacity, or physical degrees of freedom. Our results show that, under these conditions, contextual dependence cannot be represented classically without incurring an irreducible information-theoretic cost.
From this perspective, phenomena such as context sensitivity, order effects, and apparent violations of classical probabilistic reasoning need not be viewed as anomalies or implementation flaws. Instead, they arise naturally as consequences of representational constraints imposed by single-state reuse. Classical representations fail not because they are insufficiently expressive, but because they require additional contextual information that cannot be accommodated within a fixed internal state space.
More broadly, our analysis suggests that nonclassical probabilistic frameworks provide a principled means of representing contextual dependence in bounded adaptive systems. Their relevance lies not in any specific physical interpretation, but in their ability to maintain internal consistency across interventions without expanding representational resources. In this sense, contextuality reflects a fundamental trade-off between adaptability and representational economy, highlighting a general constraint on the design and analysis of intelligent systems.
VII Discussion and Outlook
We have shown that maintaining a fixed internal state space across multiple interventions entails an irreducible information-theoretic cost for classical probabilistic representations whenever contextual dependence is present. This result reframes contextuality as a representational constraint rather than a peculiarity of specific physical theories. The obstruction we identify is principle-level: it is independent of task details, dynamics, learning rules, or physical implementation.
Several clarifications are in order. First, our results do not imply that classical models are universally inadequate, nor that contextual dependence cannot be handled by classical means in general. Rather, the obstruction arises specifically under the requirement of single-state reuse without context-dependent refinement. Relaxing this requirement—by allowing explicit context encoding, additional memory, or intervention-indexed state spaces—removes the obstruction at the cost of increased representational resources. Our analysis thus makes explicit a trade-off that is often left implicit in classical modeling.
Second, while nonclassical probabilistic frameworks naturally accommodate contextual dependence under single-state constraints, our results do not mandate any particular formalism. Quantum probability appears here as one consistent realization of a broader class of nonclassical representations, distinguished not by physical assumptions but by representational structure. Identifying the minimal mathematical ingredients required for such representations remains an open direction for future work.
Finally, we emphasize that an intervention is not assumed to store its identity within the internal state, nor to leave any persistent record accessible to subsequent interventions. Operationally, interventions are memoryless transformations: once the outcome is produced, no record of the applied intervention is retained in the system. This assumption parallels standard physical idealizations such as projective measurement followed by state update, or controlled operations with subsequent reset. Allowing interventions to leave persistent traces would amount to introducing additional memory resources, thereby explicitly violating the single-state constraint.
Looking ahead, the representational perspective developed here suggests several avenues for further investigation. These include sharpening the relationship between information-theoretic cost and representational complexity, exploring extensions to dynamical and learning settings, and examining how similar obstructions manifest in multi-agent or hierarchical systems. More broadly, our results point toward a unified view of contextuality as a fundamental constraint on representation in bounded adaptive systems, bridging concepts from information theory, probability, and the theory of intelligence.
Related work on information- and memory-cost analyses of simulating contextual correlations has been developed in quantum information theory. Our contribution is not to optimize simulation costs for particular contextuality scenarios, but to isolate a representation-level obstruction under the single-state reuse requirement with interventions-as-operations rather than context labels. In this sense, the present theorem is formulated as a general representational principle, independent of specific inequality families or physical implementations.
Appendix A Proof of Theorem 1
This appendix provides a proof of the information-theoretic obstruction stated in Theorem 1. The proof is representation-theoretic in nature and does not rely on any assumptions about dynamics, learning rules, or physical realization.
Throughout this appendix, we express the abstract shorthand $H>I$ used in the main text in the explicit form
$H(M)\geq I(C;O\mid S),$
where $C$ denotes the intervention, $S$ the internal state, $O$ the observable outcome, and $M$ an auxiliary contextual variable.
A.1 Setup
We consider a classical probabilistic representation with a fixed internal state space $\mathcal{S}$, satisfying the single-state and consistency conditions of Definitions 1–3. Let $S\in\mathcal{S}$ denote the internal state random variable (with realizations $s$), and let $\mathcal{I}=\{I_{1},I_{2},\ldots\}$ denote a set of interventions. Observable outcomes are generated according to $p(o\mid s,I_{k})$.
By assumption, all observable statistics across interventions admit a single underlying classical probability space. Contextual dependence is present when no joint distribution over outcomes exists that is consistent with all observed marginals without reference to the intervention index.
A.2 Contextual Dependence and Incompatible Conditionals
Under the single-state constraint, the same internal state $s$ must support all interventions. Contextual dependence therefore manifests as incompatibility among the families of conditional distributions $\{p(o\mid s,I_{k})\}_{k}$ when viewed as arising from a single underlying random variable $s$.
To reconcile these incompatible conditionals within a classical probability space, additional information specifying how the representation should resolve contextual variation is required. This information cannot be encoded in $s$ itself without introducing intervention-dependent refinement of the state space, which is prohibited by the single-state condition.
A.3 Information-Theoretic Bound
In any classical single-state realization with an auxiliary contextual register $M$, outcomes are generated by a channel $C\to(S,M)\to O$, i.e., $p(o\mid s,m,c)=p(o\mid s,m)$ once $M$ carries all context-dependence not contained in $S$. By the data-processing inequality,
$I(C;O\mid S)\;\leq\;I(C;M\mid S).$ (2)
Moreover, $I(C;M\mid S)\leq H(M)$ holds for any random variables. Hence
$I(C;O\mid S)\;\leq\;H(M).$ (3)
Finally, contextual dependence implies $I(C;O\mid S)>0$ for any single-state classical representation that keeps $S$ context-independent; therefore $H(M)>0$ is necessary, yielding Eq. (1). The bound $H(M)\geq I(C;O\mid S)$ is saturated when $M$ carries precisely the minimal information required to distinguish incompatible conditional structures, for example when $M=f(C)$ in the construction of Sec. IV, and no additional correlations with $S$ or $O$ are introduced.
A.4 Independence from Capacity and Dynamics
Importantly, the bound above does not arise from insufficient capacity of the internal state space. Even if $\mathcal{S}$ is arbitrarily large, the requirement that it remain fixed and reused across interventions prevents the absorption of contextual information into $s$ without violating representation consistency. Likewise, the argument does not depend on any specific dynamics or learning procedures, relying solely on representational constraints.
This completes the proof.
Acknowledgements.
This work was supported by SOBIN Institute LLC under Research Grant SP005. The author used ChatGPT5.2 for English editing and takes full responsibility for the content.